A sequence is defined recursively by f(1) = 16 and f(n) = f(n-1) + 2n
Find f(4)

Answer by robertb(5830)

f(2) = f(1) + 4
f(3) = f(2) + 6
f(4) = f(3) + 8
Adding corresponding sides of the three equations above, we get
f(4) = f(1) + 18
==> f(4) = 16 + 18 = 34